OpenStudy (anonymous):

Fred is designing a rectangular poster to contain 75 in^2 of printing with a 6in margin at the top and bottom and a 2in margin at each side. What overall dimensions will minimize the amount of paper used? Can anyone help? I got stuck:/

OpenStudy (turingtest):

\[xy=75\to y=75/x\]\[(x+2)(y+6)=P(x)=(x+2)({75\over x}+6)\]find P'(x) and set it equal to zero to find x